During 2020, Sweet Company started a construction job with a contract price of $1,610,000. The job was completed in 2022. The following information is available.

2020

2021

2022

Costs incurred to date

$380,000 $847,770 $1,069,000

Estimated costs to complete

620,000 253,230 –0–

Billings to date

302,000 896,000 1,610,000

Collections to date

272,000 812,000 1,418,000

Compute the amount of gross profit to be recognized each year, assuming the percentage-of-completion method is used.

Gross profit recognized in 2020

$enter a dollar amount

Gross profit recognized in 2021

$enter a dollar amount

Gross profit recognized in 2022

$enter a dollar amount
Answer #1

Answer

  • Requirement: Percentage of completion method

2020

2021

2022

Gross Profits (Loss)

$231,800

$160,130

$149,070

  • Workings

Working

2020

2021

2022

A

Contract Price

$1,610,000

$1,610,000

$1,610,000

B

Cost Incurred to Date

$380,000

$847,770

$1,069,000

C

Estimated cost yet to be incurred to complete the contract

$620,000

$253,230

$0

D = B+C

Total Cost

$1,000,000

$1,101,000

$1,069,000

E = (B/D) x 100

% of Completion

38.00%

77.00%

100.00%

F = A x E

Revenue to date

$611,800

$1,239,700

$1,610,000

G

Revenue of Previous year

$0

$611,800

$1,239,700

H = F - G

Net Revenue this year

$611,800

$627,900

$370,300

I (=B)

Cost to date

$380,000

$847,770

$1,069,000

J

Cost to date of previous year

$0

$380,000

$847,770

K = I - J

Net Cost for the year

$380,000

$467,770

$221,230

L = H - K

Gross Profits

$231,800

$160,130

$149,070